OpenStudy (anonymous):

1.4x + 2.1 = 0.7x - 0.77

OpenStudy (anonymous):

1.4x+2.1=0.7x-0.77 Since 0.7x contains the variable to solve for, move it to the left-hand side of the equation by subtracting 0.7x from both sides. 1.4x+2.1-0.7x=-0.77 Since 1.4x and -0.7x are like terms, add -0.7x to 1.4x to get 0.7x. 0.7x+2.1=-0.77 Since 2.1 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 2.1 from both sides. 0.7x=-2.1-0.77 Subtract 0.77 from -2.1 to get -2.87. 0.7x=-2.87 Divide each term in the equation by 0.7. (0.7x)/(0.7)=-(2.87)/(0.7) Cancel the common factor of 0.7. x=-(2.87)/(0.7) Divide -2.87 by 0.7 to get -4.1. x=-4.1